OpenStudy (anonymous):

Which of these is an improper fraction? four fifths, zero ninths, eleven twelve's, five sevenths

OpenStudy (anonymous):

9/14 4/7=8/14 & 5/7=10/14 so 9/14 is in between

OpenStudy (unklerhaukus):

for your third option do you mean eleven twelfths ?

OpenStudy (anonymous):

yes I meant "twelfths" :)

OpenStudy (unklerhaukus):

proper fractions have the numerator as a number less than the denominator

OpenStudy (anonymous):

yes I remember now...Thanks!!

OpenStudy (unklerhaukus):

but all these options satisfy this rule, hmm

OpenStudy (unklerhaukus):

i think for this question the want you to see that one of these 'fractions' in the options isn't a fraction at all (after simplification)

OpenStudy (unklerhaukus):

can you see the odd one out?